Three solution of strong electrolyte 25ml of .01 M HX , 25ml of 0.1 M h2Y and 50 ml of 0.1 N Z(oh)2 are mixed what is the POH of the solution 
 

Grade:12th pass

1 Answers

SALETH
5 years ago
M1V1+M2V2+M3V3=M4V4
0.01x25)+(0.1x25)+(0.05x50) = M4V4
                                                      = 5.25 in 100ml
pH = -log[H+]
      = - log (5.25). = -0.553
pOH = 14 - pH
         = 14 - (-0.553). = 13.44